OpenStudy (anonymous):

Simplify the given expression to rational exponent form, justify each step by identifying the properties of rational exponents used. All work must be shown.

OpenStudy (anonymous):

\[1 \over \sqrt[3]{x^-6}\]

ganeshie8 (ganeshie8):

A radical can be written as rational exponent, \(\huge \frac{1}{\sqrt[3]{x^-6}} \) \(\huge \frac{1}{x^\frac{-6}{3}} \)

ganeshie8 (ganeshie8):

3 goes in 6, so simplify exponent in the bottom

ganeshie8 (ganeshie8):

\(\huge \frac{1}{x^{-2}} \)

ganeshie8 (ganeshie8):

using the negative exponent property, that simplifies to what ?

OpenStudy (anonymous):

I'm not sure :S, i'm new to exponents lol

OpenStudy (anonymous):

would you divide -2 by 1?

ganeshie8 (ganeshie8):

negative exponent property says :- \(\large \color{red}{\frac{1}{a^{-m}} = a^m} \)

OpenStudy (anonymous):

x^2?

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

alright gimme a sec to write this down xD